the width of a rectangle is 3 less than the length. the area is 10. find the dimensions of the rectangle.
Answer by checkley77(12844)

W=L-3
L(L-3)=10
L^2-3L-10=0
(L-5)(L+2)=0
L-5=0
L=5 ANSWER FOR THE LENGTH.
W=5-3
W=2 ANSWWER FOR THE WIDTH.
PROOF
2*5=10
10=10